# B-cos LM: Efficiently Transforming Pre-trained Language Models for Improved Explainability

### v3

@@ Appendix D Implementation Details
For all PLMs used in the experiments, we use the uncased base version from huggingface (Wolf et al., 2020). For both conventional models and B-cos LMs, we train them for 5 epochs with 10% linear warm-up steps on the downstream task datasets. The learning rates are set to 2e-5 for IMDB and HateXplain, and 3e-5 for AG News. All models use a batch size of 16 and a maximum sequence length of 512. For validation, we randomly sample half of the test set from IMDB and AG News.
@@
For IxG and ShapSampl, we use the Captum (Kokhlikyan et al., 2020) implementations. [^4] https://captum.ai/api/ We implement the Attention method ourselves, and LIME is sourced from the lit library [^5] https://github.com/PAIR-code/lit. For DecompX [^6] https://github.com/mohsenfayyaz/DecompX and SIG [^7] https://github.com/josephenguehard/time_interpret, we use their official implementations with default configurations. The number of samples is set to 25 for ShapSampl and 3,000 for LIME, with [MASK] as the baseline token. For all explanation methods at the embedding level, model predictions are attributed to the combined sum of word, position, and token type embeddings (if applicable). In the main experiments, we compute token attribution scores by summing over all embedding dimensions, as this approach demonstrates better faithfulness results than using the L2 norm.
@@
In Section 7, we follow Ferrando et al. (2023) to generate contrastive explanations that highlight why the models predicts the target token instead of the foil token. For Occlusion explanations, we use the [PAD] token to perform occlusion, instead of a zero vector as done in Yin & Neubig (2022) and Ferrando et al. (2023). Using zero vectors distorts the input distribution and, in generative settings, can influence predictions differently depending on position. To avoid such positional effects, we instead occlude using the in-distribution embedding of the [PAD] token.
@@
We follow Ferrando et al. (2023) to use the following nine BLiMP subsets with corresponding IDs. aga: anaphor_gender_agreement; ana: anaphor_number_agreement; asp: animate_subject_passive; dna: determiner_noun_agreement_1; dnai: determiner_noun_agreement_irregular_1; dnaa: determiner_noun_agreement_with_adj_1; dnaai: determiner_noun_agreement_with_adj_irregular_1; npi: npi_present_1; darn: distractor_agreement_relational_noun. Examples of these datasets and the IOI dataset can be found in Table 7.
@@ Appendix F SeqPG Example
Figure 7 presents a SeqPG example from AG News using B-cos BERT. For better visualization, each segment is truncated to 20 tokens here instead of 50 used in the experiments. Unlike the hybrid document evaluation proposed by Poerner et al. (2018), our approach explicitly controls segment length and position to ensure a fair comparison. Additionally, we measure the proportion of correctly assigned positive attributions rather than relying solely on the highest attribution value.
@@ Appendix G Task Performance of Other B-cos LMs
Figure 8: Mean accuracy of conventionally fine-tuned and B-cos DistilBERT models averaged over three runs. We use B=1.5, 1.25, and 1.5 for AG News, IMDB and HateXplain, respectively. B-cos models perform comparably to conventional models on most tasks.
Figure 9: Mean accuracy of conventionally fine-tuned and B-cos RoBERTa models averaged over three runs. We use B=1.5, 1.25, and 1.5 for AG News, IMDB and HateXplain, respectively. B-cos models perform comparably to conventional models on most tasks.
@@ Appendix H Faithfulness Evaluation of Other B-cos LMs
Table 9: Faithfulness evaluation for conventionally fine-tuned DistilBERT and B-cos DistilBERT across three datasets. We use B=1.5, 1.25, and 1.5 for AG News, IMDB and HateXplain, respectively. The best results are in bold. We find that B-cos explanations are consistently more faithful than post-hoc explanations from both models.
@@
Table 10: Faithfulness evaluation for conventionally fine-tuned RoBERTa, Saloss RoBERTa and B-cos RoBERTa across three datasets. We use B=1.5, 1.25, and 1.5 for AG News, IMDB and HateXplain, respectively. The best results are in bold. We find that B-cos explanations are consistently more faithful than post-hoc explanations from both models.
@@ Appendix I Comparison to Rationale-Based Models
We compare B-cos LMs to one rationale-based, explain-then-predict BERT model, RGFS-SA (Saha et al., 2023) [^10] https://huggingface.co/Hate-speech-CNERG/Rationale_predictor on HateXplain. This model leverages human rationales as additional supervision during training. As shown in Table 11, although the RGFS-SA model brings improvement over the conventional BERT model, it generates considerably less faithful rationales compared to B-cos explanations.
@@
Figure 13 presents two examples where B-cos explanations do not align with human expectations. As B-cos explanations faithfully summarize model computations, they may also reveal spurious correlations from the training data and occasionally produce unreasonable rationales. For example, in the first case, the token “[user]” receives high attribution, likely due to recurring patterns in the data. Another observed error is polarity misinterpretation: in the second case, the token “niggas” is incorrectly highlighted as positive evidence for a non-toxic prediction. Similar types of undesired behaviors have also been reported for other explanation methods.
@@
Overall, larger B values generally lead to stronger input-weight alignment compared to smaller B and conventional models, as evidenced by the curves for B=1.5 and B=2.5 lying above those for the conventional model and B=1. However, the alignment pattern becomes more complex when comparing B=1.5 and B=2.5. Specifically, at B=2.5, the most aligned input-weight pairs exhibit higher alignment than in other models, but some pairs show very low alignment. This result may arise because certain weights are highly optimized for specific input patterns, leading to poor alignment with others, particularly in later layers where input features become more anisotropic (Ethayarajh, 2019; Li et al., 2020). As a result, some outputs from the B-cos layers are highly negative. When these outputs are fed into GELU activation functions, their dynamic weights approach zero, making the explanations more sparse.
@@
Table 12: Task performance, explanation faithfulness, and explanation entropy of B-cos BERT models on HateXplain with different B values. Results are averaged over three runs. Similar to Figure 5, task performance and explanation faithfulness peak around B=1.5, while explanation entropy negatively correlates with B.
@@
Dataset Random Grad Norm IxG Occlusion Logit ALTI Logit B-cos
@@
Table 14: MRR Alignment of different explanation methods on GPT-2 small predictions on every dataset. B-cos explanations are extracted from the B-cos GPT-2 model. Logit and ALTI Logit results are duplicated from Ferrando et al. (2023).
Dataset Random Grad Norm IxG Occlusion B-cos
@@
Table 15: MRR Alignment of different explanation methods on Llama-3.2 predictions on every dataset. B-cos explanations are extracted from the B-cos Llama-3.2 model. As Llama models are not supported in Ferrando et al. (2023), we do not include their results.

### v4

@@ Appendix D Implementation Details
For all PLMs used in the experiments, we use the uncased base version from huggingface (wolf-etal-2020-transformers). For both conventional models and B-cos LMs, we train them for 5 epochs with 10% linear warm-up steps on the downstream task datasets. The learning rates are set to 2e-5 for IMDB and HateXplain, and 3e-5 for AG News. All models use a batch size of 16 and a maximum sequence length of 512. For validation, we randomly sample half of the test set from IMDB and AG News.
@@
For IxG and ShapSampl, we use the Captum (DBLP:journals/corr/abs-2009-07896) implementations. [^4] https://captum.ai/api/ We implement the Attention method ourselves, and LIME is sourced from the lit library [^5] https://github.com/PAIR-code/lit. For DecompX [^6] https://github.com/mohsenfayyaz/DecompX and SIG [^7] https://github.com/josephenguehard/time_interpret, we use their official implementations with default configurations. The number of samples is set to 25 for ShapSampl and 3,000 for LIME, with [MASK] as the baseline token. For all explanation methods at the embedding level, model predictions are attributed to the combined sum of word, position, and token type embeddings (if applicable). In the main experiments, we compute token attribution scores by summing over all embedding dimensions, as this approach demonstrates better faithfulness results than using the L2 norm.
@@
In Section 7, we follow ferrando-etal-2023-explaining to generate contrastive explanations that highlight why the models predicts the target token instead of the foil token. For Occlusion explanations, we use the [PAD] token to perform occlusion, instead of a zero vector as done in yin-neubig-2022-interpreting and ferrando-etal-2023-explaining. Using zero vectors distorts the input distribution and, in generative settings, can influence predictions differently depending on position. To avoid such positional effects, we instead occlude using the in-distribution embedding of the [PAD] token.
@@
We follow ferrando-etal-2023-explaining to use the following nine BLiMP subsets with corresponding IDs. aga: anaphor gender agreement; ana: anaphor number agreement; asp: animate subject passive; dna: determiner noun agreement 1; dnai: determiner noun agreement irregular 1; dnaa: determiner noun agreement with adj 1; dnaai: determiner noun agreement with adj irregular 1; npi: npi present 1; darn: distractor agreement relational noun. Examples of these datasets and the IOI dataset can be found in Table 7.
@@ Appendix F SeqPG Example
Figure 7 presents a SeqPG example from AG News using B-cos BERT. For better visualization, each segment is truncated to 20 tokens here instead of 50 used in the experiments. Unlike the hybrid document evaluation proposed by poerner-etal-2018-evaluating, our approach explicitly controls segment length and position to ensure a fair comparison. Additionally, we measure the proportion of correctly assigned positive attributions rather than relying solely on the highest attribution value.
@@ Appendix G Task Performance of Other B-cos LMs
Figure 8: Mean accuracy of conventionally fine-tuned and B-cos DistilBERT models averaged over three runs. B-cos models perform comparably to conventional models on most tasks.
Figure 9: Mean accuracy of conventionally fine-tuned and B-cos RoBERTa models averaged over three runs. B-cos models perform comparably to conventional models on most tasks.
@@ Appendix H Faithfulness Evaluation of Other B-cos LMs
Table 9: Faithfulness evaluation for conventionally fine-tuned DistilBERT and B-cos DistilBERT across three datasets. The best results are in bold. We find that B-cos explanations are consistently more faithful than post-hoc explanations from both models.
@@
Table 10: Faithfulness evaluation for conventionally fine-tuned RoBERTa, Saloss RoBERTa and B-cos RoBERTa across three datasets. The best results are in bold. We find that B-cos explanations are consistently more faithful than post-hoc explanations from both models.
@@ Appendix I Comparison to Rationale-Based Models
We compare B-cos LMs to one rationale-based, explain-then-predict BERT model, RGFS-SA (punyajoy-etal-2023-rationale) [^10] https://huggingface.co/Hate-speech-CNERG/Rationale_predictor on HateXplain. This model leverages human rationales as additional supervision during training. As shown in Table 11, although the RGFS-SA model brings improvement over the conventional BERT model, it generates considerably less faithful rationales compared to B-cos explanations.
@@
Figure 13 presents two examples where B-cos explanations do not align with human expectations. As B-cos explanations faithfully summarize model computations, they may also reveal spurious correlations from the training data and occasionally produce unreasonable rationales. For example, in the first case, the token "[user]" receives high attribution, likely due to recurring patterns in the data. Another observed error is polarity misinterpretation: in the second case, the token "niggas" is incorrectly highlighted as positive evidence for a non-toxic prediction. Similar types of undesired behavior have also been reported for other explanation methods.
@@
Overall, larger B values generally lead to stronger input-weight alignment compared to smaller B and conventional models, as evidenced by the curves for B=1.5 and B=2.5 lying above those for the conventional model and B=1. However, the alignment pattern becomes more complex when comparing B=1.5 and B=2.5. Specifically, at B=2.5, the most aligned input-weight pairs exhibit higher alignment than in other models, but some pairs show very low alignment. This result may arise because certain weights are highly optimized for specific input patterns, leading to poor alignment with others, particularly in later layers where input features become more anisotropic (ethayarajh-2019-contextual; li-etal-2020-sentence). As a result, some outputs from the B-cos layers are highly negative. When these outputs are fed into GELU activation functions, their dynamic weights approach zero, making the explanations more sparse.
@@
Table 12: Task performance, explanation faithfulness, and explanation entropy of B-cos BERT models on HateXplain with different B values. Results are averaged over three runs. Similar to Figure 5, task performance and explanation faithfulness peak around B=1.5, while explanation entropy correlates negatively with B.
@@
Dataset Random Grad Norm IxG Occlusion Logit ALTI Logit B-cos GPT-2
@@
Table 14: MRR Alignment of different explanation methods on GPT-2 small predictions on every dataset. B-cos explanations are extracted from the B-cos GPT-2 model. Logit and ALTI Logit results are duplicated from ferrando-etal-2023-explaining.
Dataset Random Grad Norm IxG Occlusion B-cos Llama-3.2
@@
Table 15: MRR Alignment of different explanation methods on Llama-3.2 predictions on every dataset. B-cos explanations are extracted from the B-cos Llama-3.2 model. As Llama models are not supported in ferrando-etal-2023-explaining, we do not include their results.